# Maximal Extractable Value in Decentralized Finance: Taxonomy, Detection, and Mitigation

### v1

@@ 7 MEV Mitigation Strategies
To mitigate the negative impacts of MEV, whether by reducing its occurrence or democratizing its exploitation, various strategies have been developed. These strategies focus on designing fair transaction ordering policies, executing privacy-preserving transactions, or developing protocols for processing transactions using public pools. In this section, we explore different categories of MEV mitigation strategies, evaluating their effectiveness in addressing different types of MEV while maintaining decentralization and security.
@@ 7.1 Transaction Ordering
The most well-known strategy based on submission or arrival times is order fairness mechanisms such as receive order fairness, send order fairness, approximate order fairness, and block order fairness [32]. Receive order fairness mechanism states that if a majority (more than 50%) of the validator nodes receives transaction before , then should precede in the final order of transactions [38]. On the other hand, send order fairness states that if transaction is sent before , then should precede in the final order of transactions. Intel SGX Trusted Execution Environment (TEE) [39] or ARM Trustzone [40] could be used to verify and trust the timestamps of these transactions [41]. However, due to network delays, achieving exact transaction ordering under the sender order fairness mechanism may not always be feasible [41]. Consequently, approximate order fairness allows for flexibility, relaxing the strict sequencing of transactions and if validators receive them within a specified time window. If the majority of validator nodes receive at least rounds earlier than , then all honest validator nodes should not deliver until has been delivered [42]. These order fairness mechanisms primarily consider front-running, back-running, sandwich, arbitrage, and liquidation MEV transactions by focusing on transaction-level ordering. To address time-bandit MEV, a block order fairness mechanism is proposed. When transactions and are received by the validators within a specified time window, block order fairness mechanism ensures that the majority of validators receive before , then all honest validator nodes should not deliver at a later block index than it delivers [38, 42]. Temporal order fairness strategies can be further extended by discretizing the time into periods measured in terms of time units or the number of blocks [19]. If multiple searchers submit competing transactions, then the dispute can be resolved using the Solomonic approach [43]. This approach assumes that each transaction initiator (whether searcher or legitimate user) knows its own status. An initiator (selected randomly) is given the opportunity to withdraw their transaction. If the transaction is withdrawn, the other initiator’s transaction is executed, else the transaction amount is burned. If the legitimate initiator is given the opportunity to withdraw, they incur a loss. In contrast, if a searcher withdraws, they do not incur a loss. Both initiators incur transaction fees if they assert their transactions, which discourages searchers from making illegitimate claims. However, this approach is only suitable for mitigating replacement front-running. Furthermore, implementing and managing the random selection and token-burning mechanisms could be complex. In addition, the strategy becomes an iterative process with an increasing number of competing searchers, leading to increased costs and higher latency.
@@
The order fairness [32, 19] strategies rely solely on transaction timestamps, which could create an unfair advantage for searchers with low-latency infrastructure [8]. To address this limitation, Mamageishvili et al. [33] proposed TimeBoost, an ordering strategy based on both transaction timestamps and gas fees. TimeBoost calculates a score for each transaction received within a specific time interval and orders transactions in descending order based on these scores. The score is determined by subtracting the transaction arrival time from a bidding function, which is defined as the ratio of and . The transaction priority increases with the bid. Consequently, searchers may compete by resubmitting transactions with progressively increasing bids, potentially leading to network congestion and instability. Furthermore, TimeBoost could disproportionately favor wealthy searchers who can invest in low-latency infrastructure and afford to place high bids.
@@
To prevent favoring wealthy searchers, transaction ordering strategies based on random [34] or deterministic [35] sequencers have been proposed. Sinai and Hoh In [34] proposed the Quantum Random Transaction Ordering Protocol (Q-RTOP), which utilizes a quantum random number generator to finalize the order of transactions within a block. Empirical results demonstrate the effectiveness of Q-RTOP in randomizing over 8000 transactions in under 25 milliseconds. However, Q-RTOP does not guarantee that block producers will refrain from rearranging the randomized transactions to extract MEV, as proving that a block contains truly randomized transactions remains challenging. On the other hand, Doe et al. [35] proposed a deterministic sequencer strategy for transaction ordering based on transaction confidentiality and delay tolerance. The authors introduced a Weighted Sequencing Service (WSS) that computes the weight for each transaction based on delay tolerance, confidentiality, and transaction workload. Transactions are then ordered based on their weights. However, searchers may manipulate delay tolerance and confidentiality requirements of their transactions to influence the ordering.
@@
In the context of transaction ordering based on MM designs, Ciampi et al. [36] proposed Fair Market Maker (FairMM), an AMM based on the protocol that utilizes off-chain communication for transaction ordering. In this protocol, a buyer creates a smart contract specifying the required token amount and locks it on the blockchain. The buyer then sends their identity (without token details) to the seller off-chain, and the seller responds with the exchange rate. If the buyer agrees, the contract initiates the trade by forwarding trade details to a MM algorithm. The MM can then accept or decline to trade. If the MM conducts trades with other traders in between, a transaction reordering will be flagged. While FairMM does not explicitly prevent transaction reordering, it discourages malicious activities by publicly flagging any misbehavior. Furthermore, it cannot mitigate back-running transactions since a searcher can execute a transaction that back-runs the victim transaction without requiring any reordering. Empirical analysis shows that, compared to the popular AMM Uniswap, FairMM results in constant gas costs, higher throughput, and faster trade execution. Similarly, Moosavi and Clark [37] proposed Lissy, an alternative to traditional AMM for fair, efficient, and transparent token trading. Unlike continuous trading in an AMM, Lissy employs a call market mechanism for periodic trading, collecting orders over a set period and executing them simultaneously at a single clearing price. This approach eliminates front-running, where traders exploit timing differences to gain an advantage. However, Lissy faces scalability issues in the main Ethereum network due to periodic execution.
@@ 7.2 Privacy-preserving Public Pools
Kavousi et al. [44] extended the threshold decryption method by proposing BlindPerm, a framework designed to enhance security against MEV transaction reordering. BlindPerm shuffles the order of committed transactions using a seed generated through a permutation technique. This added layer of permutation ensures that even after the transactions are decrypted, their order remains concealed, providing additional protection against MEV exploits. A similar approach is employed by Piet et al. [45], where encrypted transactions are randomly shuffled using a randomness beacon. This beacon utilizes Ethereum’s RANDAO algorithm, which generates a random value by combining shares of input from a committee. The input is generally a digital signature of known information. However, threshold decryption introduces high bandwidth overheads because each committee member must propagate their share of the decrypted transaction for every encrypted transaction in the network. This increases the amount of data that needs to be transmitted and processed, leading to network congestion and scalability constraints. Considering an example scenario of 500 transactions, each of size 64 bytes, and a committee of 100 members, the additional data load on the network due to threshold decryption would be lower-bounded at 1632 KB if the threshold is set at 51%, and upper-bounded at 3200 KB if the threshold is set at 100%.
To reduce the bandwidth overheads associated with threshold decryption, Momeni et al. [18] proposed FairBlock, which uses a global public key and a block identifier (for instance, block index) to encrypt transactions. During the decryption phase, FairBlock generates a shared key for the entire block based on the block identifier, rather than requiring a shared key for each individual transaction as in threshold decryption. This approach significantly reduces bandwidth overheads and improves scalability. Experimental results show that the message size of FairBlock is between 0.4% and 5.26% of the message size required by threshold decryption approaches. However, FairBlock introduces a risk of a single point of failure. This is because if the shared key is compromised, all transactions within the blocks could be exposed to MEV exploitation.
@@ 7.3 Private Pools
To democratize MEV, private pool mitigation strategies focus on privately submitting transactions to block producers rather than the public mempool. The Flashbots project181818https://www.flashbots.net/, accessed on 05 September 2024 introduced the concept of private pools and relays. The searchers submit their transactions to relays off-chain, avoiding public mempool exposure. Relays collect these transactions in a private mempool, where the gas fee bids remain sealed, unlike the open-bid process in PGAs. The relays then forward the transactions to block producers. Block producers pack the transactions (both from public mempool and forwarded by the relays) into blocks, taking into account the block size limit. Transactions are ordered by decreasing gas prices within the block. By replacing the open-bid on-chain auction with a sealed-bid off-chain auction, Flashbots effectively mitigate the issues of network congestion and higher gas prices associated with PGAs [6].
@@
The first Flashbots block was mined at block height 11,834,049 on February 11, 2021, after the project launched earlier that year [28]. Experimental results demonstrated that the sealed-bid auction approach results in lower average bidding prices compared to the conventional open-bid mechanism, contributing to more stable gas prices [47]. However, Weintraub et al. [28] argue that the Flashbots project may not optimally democratize and transparently distribute MEV. Their findings suggest that the benefits of Flashbots are disproportionately skewed toward block producers, with searchers sometimes experiencing negative profits, thus failing to achieve fair profit distribution.
@@
After Ethereum’s transition from PoW to PoS, an architectural change introduced Proposer-Builder Separation (PBS) [6, 8], dividing the block producer role into two distinct roles performed by block builder and block proposer. In PBS, searchers submit transactions privately to block builders. They can also submit a group of ordered transactions, called bundles, that should be mined together in a specific order to ensure they capture MEV. Transactions from users other than searchers can also be submitted privately to the block builder. The block builder creates a block from both public and private transactions, aiming to maximize gas fees. The block is then sent to relays, which act as intermediaries between the block builder and the block proposer. The relay validates the block and forwards only the block header information (excluding transaction details) to the proposer. The block proposer selects the most profitable block (based on block rewards) from those received from multiple relays and returns the signed block to the chosen relay. The proposer then receives the block contents and broadcasts them to the network. Searchers may entice builders to include their bundles in a block by offering higher transaction fees, sharing a portion of the MEV, and/or providing private incentives. In turn, builders might encourage proposers to choose their blocks by offering a share of transaction fees/MEV obtained from searchers and/or by providing other private incentives. However, PBS suffers from the risk of centralization, as a builder relay collusion could indulge in block censoring or MEV exploitation activities [48].
@@
Table IV summarizes the effectiveness of mitigation strategies in addressing different types of MEV. Furthermore, it discusses the limitations of each strategy, providing guidance for developers and policy-makers in designing and implementing optimal MEV mitigation strategies.
@@ 8 MEV Simulation and Extraction Methods
In the context of MEV simulation, Qin et al. [15] proposed a method to simulate replacement front-running by duplicating a victim transaction, replacing only the sender’s address by the searcher’s address. The transaction is then executed locally on the highest block. If profitable, the searcher front-runs the victim transaction. A similar approach was developed by Stucke et al. [51] to simulate displacement front-running transactions. The approach executes the transactions locally on a Ganache fork to compute profitability. The approach also simulates sandwich transactions. This is by identifying victim transactions that perform swaps with a liquidity pool. It then front-runs the victim transaction by performing the same swap and back-runs it a reverse swap. Bogatyy and his team developed a simulator to perform displacement front-running transactions on the Bancor platform 191919https://github.com/bogatyy/bancor, accessed on 05 September 2024. The simulator monitors the mempool for a victim transaction and executes a displacement transaction on the local Ethereum node. If profitable, the transaction is submitted to the network with a gas price higher than that of the victim via JSON RPC. The Subway bot 202020https://github.com/libevm/subway, accessed on 05 September 2024 simulates sandwich transactions by monitoring the mempool for profitable opportunities. It determines optimal swap values and gas prices for front-running and back-running transactions, then executes them to capture profit.
@@
Regarding MEV extractions, Schwarz-Schilling et al. [52] and Öz et al. [53] discovered that the bid value submitted by a block builder in PBS increases with bid submission time. This increase might be attributed to the expansion of the public mempool and/or the submission of more private transaction bundles to the proposer. Consequently, the authors suggested that an honest but rational block proposer might delay signing a block to optimize MEV. However, empirical analysis in these studies showed that proposers do not engage in such waiting games, likely due to concerns about potential reputation damage associated with self-serving behaviors. To optimize MEV, Babel and Baker [54] proposed MEV-peri, a peer selection approach that select peers based on the value of transactions relayed by the peers and the latency of the received transactions. Empirical results reveal that peer selection using MEV-peri extracts more MEV compared to random and latency-based peer selection approaches.
@@
Bartoletti et al. [22] proposed a method to optimize MEV from dagwood sandwich transactions by modeling a transaction as a single-player single-round game. Similarly, Wang et al. [24] proposed a method to optimize MEV in the case of burger sandwich transactions. The proposed method involves a transaction selecting algorithm and an optimal attack algorithm. The selecting algorithm uses an iterative approach to determine a set of victim transactions that would yield optimal MEV. The optimal attack algorithm employs binary search to output parameters for an optimal front-running transaction based on the set of victim transactions. Babel et al. [50] introduced Lanturn, a learning-based framework for simulating MEV optimization strategies. It consists of an optimizer and a simulation environment. The optimizer, inspired by Genetic Algorithms, uses a bi-loop approach to develop a transaction sequence that maximizes profit. The outer loop determines the optimal order of transactions, while the inner loop fine-tunes transaction variables (e.g., token amounts in a swap transaction). The simulation environment evaluates the developed sequence by computing MEV at the current block height, executing transactions in a FIFO (First In, First Out) order on the local node. Experimental results showed that Lanturn extracts more MEV than Flashbots. Zhou et al. [55] introduced DEFIPOSER-ARB and DEFIPOSER-SMT to generate cyclic and non-cyclic arbitrage transactions for MEV extraction. DEFIPOSER-ARB identifies arbitrage opportunities across different DEXes by first constructing a graph of the current blockchain state and detecting cyclic transactions with potential profit. A local search then evaluates the profit from these opportunities, and the most profitable transaction is selected for inclusion in the next block. On the other hand, DEFIPOSER-SMT exploits non-cyclic profit-generating transactions by creating a state transition model of the DeFi system. The paths in the transition model represent potential profitable transaction sequences. Heuristics are then employed to prune the search space and execute the most profitable path.

### v2

@@ 7 MEV Mitigation Strategies
To mitigate the negative impacts of MEV, whether by reducing its occurrence or democratizing its exploitation, various strategies have been developed. These strategies focus on designing fair transaction ordering policies, executing privacy-preserving transactions, or developing protocols for processing transactions using private pools. In this section, we explore different categories of MEV mitigation strategies, evaluating their effectiveness in addressing different types of MEV while maintaining decentralization and security.
@@ 7.1 Transaction Ordering
The most well-known strategy based on submission or arrival times is order fairness mechanisms such as receive, send, approximate, and block order fairness [39]. Receive order fairness states that if a majority (more than 50%) of validators receive transaction before , then should precede in the final order [45]. Send order fairness states that if is sent before , then should precede in the final order. Intel SGX Trusted Execution Environment (TEE) [46] or ARM Trustzone [47] could be used to verify and trust the timestamps of these transactions [48]. However, due to network delays, achieving strict send order fairness is often infeasible [48]. To address this, approximate order fairness relaxes strict sequencing of transactions and if validators receive them within a specified time window. If the majority of validator nodes receive at least rounds earlier than , then all honest validator nodes should not deliver until has been delivered [49]. These order fairness mechanisms primarily consider front-running, back-running, sandwich, arbitrage, and liquidation MEV transactions by focusing on transaction-level ordering. To address time-bandit MEV, block order fairness ensures that if and arrive within a time window and a majority of validators receive first, then should not be placed in a later block than [45, 49]. Temporal order fairness strategies can be further extended by discretizing the time into fixed periods measured in time units or block numbers [22]. If multiple searchers submit competing transactions, then disputes can be resolved using the Solomonic approach212121https://github.com/solomonic-mechanism, accessed on 25 February 2025 [50]. This approach assumes that each transaction initiator (whether searcher or legitimate user) knows its own status. An initiator (selected randomly) is given the opportunity to withdraw their transaction. If the transaction is withdrawn, the other initiator’s transaction is executed, else the transaction amount is burned. A Legitimate initiator incurs a loss if forced to withdraw, while a searcher does not, discouraging illegitimate claims. However, this approach is only suitable for mitigating replacement front-running. Furthermore, implementing and managing the random selection and token-burning mechanisms can be complex, and as the competition among searchers increases, costs and latency also rise.
@@
The order fairness [39, 22] strategies rely solely on transaction timestamps, potentially giving an unfair advantage to searchers with low-latency infrastructure [9]. To address this, Mamageishvili et al. [40] proposed TimeBoost, an ordering strategy based on both transaction timestamps and gas fees. TimeBoost assigns a score for each transaction received within a specific time interval and orders them in descending order. The score is calculated by subtracting the transaction arrival time from a bidding function that considers both bid amounts and total bids in the system. Higher bid increases transaction priority, prompting searchers to compete by resubmitting transactions with progressively increasing bids, which may cause network congestion and instability. Furthermore, TimeBoost could disproportionately favor wealthy searchers who can invest in low-latency infrastructure and afford to place high bids.
@@
To prevent favoring wealthy searchers, transaction ordering strategies based on random [41] or deterministic [42] sequencers have been proposed. Sinai and Hoh In [41] proposed the Quantum Random Transaction Ordering Protocol (Q-RTOP), which utilizes a quantum random number generator to finalize transaction order. However, Q-RTOP cannot prevent block producers from reordering the randomized transactions, as verifying true randomness remains challenging. On the other hand, Doe et al. [42] proposed a deterministic sequencer strategy for transaction ordering based on transaction confidentiality and delay tolerance. The authors introduced a Weighted Sequencing Service (WSS) that computes the weight for each transaction based on delay tolerance, confidentiality, and transaction workload. Transactions are then ordered based on their weights. However, searchers may manipulate delay tolerance and confidentiality requirements of their transactions to influence the ordering.
@@
In the context of transaction ordering in AMM designs, Ciampi et al. [43] introduced Fair Market Maker (FairMM), an AMM built on their proposed -Trade protocol. FairMM utilizes off-chain communication for transaction ordering. In -Trade protocol, a buyer creates a smart contract specifying the required token amount and locks it on the blockchain. The buyer then sends their identity (without token details) to the seller off-chain, and the seller responds with the exchange rate. If the buyer agrees, the contract initiates the trade by forwarding trade details to a MM algorithm. The MM can then accept or decline to trade. If the MM conducts trades with other traders in between, a transaction reordering will be flagged. While FairMM does not explicitly prevent transaction reordering, it discourages malicious activities by publicly flagging any misbehavior. Furthermore, it cannot mitigate back-running transactions since a searcher can execute a transaction that back-runs the victim transaction without requiring any reordering. Empirical analysis shows that, compared to the popular AMM Uniswap, FairMM results in constant gas costs, higher throughput, and faster trade execution. Similarly, Moosavi and Clark [44] proposed Lissy222222https://github.com/MadibaGroup/2020-Orderbook, accessed on 25 February 2025, an alternative to traditional AMM for fair, efficient, and transparent token trading. Unlike continuous trading in an AMM, Lissy employs a call market mechanism for periodic trading, collecting orders over a set period and executing them simultaneously at a single clearing price, referred to as batch auctions [51]. This approach eliminates front-running, where traders exploit timing differences to gain an advantage. However, Lissy faces scalability issues in the main Ethereum network due to periodic execution.
@@ 7.2 Privacy-preserving Public Pools
Kavousi et al. [52] extended the threshold decryption method by proposing BlindPerm, a framework designed to enhance security against MEV transaction reordering. BlindPerm shuffles the order of committed transactions using a seed generated through a permutation technique. This added layer of permutation ensures that even after the transactions are decrypted, their order remains concealed, providing additional protection against MEV exploits. A similar approach is employed by Piet et al. [53], where encrypted transactions are randomly shuffled using a randomness beacon232323https://github.com/sanjayss34/geth-random-order, accessed on 25 February 2025. This beacon utilizes Ethereum’s RANDAO algorithm, which generates a random value by combining shares of input from a committee. The input is generally a digital signature of known information. However, some threshold decryption approaches might allow the decryption of pending transactions that are not included in a block. This can expose transaction information, potentially enabling MEV extraction if a searcher is part of the committee of decryptors. Furthermore, threshold decryption introduces high bandwidth overheads because each committee member must propagate their share of the decrypted transaction for every encrypted transaction in the network. This increases the amount of data that needs to be transmitted and processed, leading to network congestion and scalability constraints. Considering an example scenario of 500 transactions, each of size 64 bytes, and a committee of 100 members, the additional data load on the network due to threshold decryption would be lower-bounded at 1632 KB if the threshold is set at 51%, and upper-bounded at 3200 KB if the threshold is set at 100%.
To reduce the bandwidth overheads associated with threshold decryption, Momeni et al. [21] proposed FairBlock242424https://github.com/fairblock, accessed on 25 February 2025, which uses a global public key and a block identifier (for instance, block index) to encrypt transactions. During the decryption phase, FairBlock generates a shared key for the entire block based on the block identifier, rather than requiring a shared key for each individual transaction as in threshold decryption. This approach significantly reduces bandwidth overheads and improves scalability. Experimental results show that the message size of FairBlock is between 0.4% and 5.26% of the message size required by threshold decryption approaches. However, FairBlock introduces a risk of a single point of failure. This is because if the shared key is compromised, all transactions within the blocks could be exposed to MEV exploitation.
@@ 7.3 Private Pools
To democratize MEV, private pool mitigation strategies focus on privately submitting transactions to block producers rather than the public mempool. The Flashbots project252525https://www.flashbots.net/, accessed on 05 September 2024 introduced the concept of private pools and relays. The searchers submit their transactions to relays off-chain, avoiding public mempool exposure. Relays collect these transactions in a private mempool, where the gas fee bids remain sealed, unlike the open-bid process in PGAs. The relays then forward the transactions to block producers. Block producers pack the transactions (both from public mempool and forwarded by the relays) into blocks, taking into account the block size limit. Transactions are ordered by decreasing gas prices within the block. By replacing the open-bid on-chain auction with a sealed-bid off-chain auction, Flashbots effectively mitigate the issues of network congestion and higher gas prices associated with PGAs [7].
@@
The first Flashbots block was mined at block height 11,834,049 on February 11, 2021, after the project launched earlier that year [35]. Experimental results demonstrated that the sealed-bid auction approach results in lower average bidding prices compared to the conventional open-bid mechanism, contributing to more stable gas prices [55]. However, the reliance on centralized relays within Flashbots raises trust concerns [56], and there is a lack of accountability regarding potential leakage of builders’ private transactions [57]. Furthermore, Weintraub et al. [35] argue that the Flashbots project may not optimally democratize and transparently distribute MEV. Their findings suggest that the benefits of Flashbots are disproportionately skewed toward block producers, with searchers sometimes experiencing negative profits, thus failing to achieve fair profit distribution.
@@
After Ethereum’s transition from PoW to PoS, an architectural change introduced PBS [7, 9], dividing the block producer role into two distinct roles performed by block builder and block proposer. In PBS, searchers submit transactions privately to block builders. They can also submit a group of ordered transactions, called bundles, that should be mined together in a specific order to ensure they capture MEV. Transactions from users other than searchers can also be submitted privately to the block builder. The block builder creates a block from both public and private transactions, aiming to maximize gas fees. The block is then sent to relays, which act as intermediaries between the block builder and the block proposer. The relay validates the block and forwards only the block header information (excluding transaction details) to the proposer. The block proposer selects the most profitable block (based on block rewards) from those received from multiple relays and returns the signed block to the chosen relay. The proposer then receives the block contents and broadcasts them to the network. Searchers may entice builders to include their bundles in a block by offering higher transaction fees, sharing a portion of the MEV, and/or providing private incentives. In turn, builders might encourage proposers to choose their blocks by offering a share of transaction fees/MEV obtained from searchers and/or by providing other private incentives. However, PBS suffers from the risk of centralization, as a builder relay collusion could indulge in block censoring or MEV exploitation activities [58]. Furthermore, although PBS benefits the participating actors, it has exacerbated the impact of MEV on regular network users. To address this, Babel et al. [59] proposed PROF (Protected Order Flow), which introduces a sequencer and bundle merger to mitigate harmful MEV while ensuring profitability for PBS participants. The sequencer collects private transactions from users and orders them into a bundle. The bundle merger, operating at PBS relays, appends this bundle to the most profitable block from builders. The proposer includes the PROF-enriched block for finalization. The authors further introduced PROF-Share, which redistributes MEV profits from back-running back to users. However, in congestion scenarios where the builder-generated bundles occupy the full block capacity, the PROF bundle may be excluded, leading to transaction delays or failures if it does not offer competitive fees to builders and proposers to include the PROF bundle.
Previously discussed mitigation strategies reduce the risk of MEV by transaction ordering. In contrast, intent-centric approaches focus on MEV-resistant transaction generation. For instance, CoWSwap262626https://cow.fi/cow-protocol, accessed on 14 February 2025 allows users to submit intents rather than submitting detailed transaction information. Third-party entities, referred to as solvers, then compete to execute the intents in batches. However, intent-centric approaches face a trade-off between privacy and efficiency. Similar to the potential collusion between relays and builders in PBS, solver collusion could lead to centralization, posing a risk to the system’s fairness.
@@
Table III summarizes the effectiveness of mitigation strategies in addressing different types of MEV. Furthermore, it discusses the limitations of each strategy, providing guidance for developers and policy-makers in designing and implementing optimal MEV mitigation strategies. While these strategies focus on reducing MEV exploitation, alternate approaches modify the underlying economic conditions that create MEV opportunities or develop MEV-resistant DeFi ecosystems. For example, Miqado [61] mitigates liquidation MEV in lending and borrowing protocols by replacing the fixed spread liquidation mechanism with a supporter-based collateral top-up approach. In traditional DeFi lending protocols, searchers exploit liquidation opportunities, but Miqado delays or mitigates this by allowing external supporters to add collateral to prevent liquidation. The supporter can later get repaid with interest by the borrower or claim the position. However, this approach depends on the availability of supporters, which in turn relies on proper incentive mechanisms. Furthermore, formal verification techniques enable the development of MEV-resistant DeFi protocols. Frameworks such as Clockwork Finance [3] aid in the development of MEV-resistant DeFi protocols by identifying potential security risks and quantifying extractable value. Similarly, KEVM enables symbolic execution for reasoning about transaction ordering and gas costs, ensuring that smart contracts are formally verified against MEV attack vectors [62].
After discussing MEV mitigation strategies, it is crucial to understand how different types of MEV transactions can be simulated, extracted, and optimized in practice. MEV simulation helps model the behavior of various MEV types and assess the effectiveness of detection approaches and mitigation strategies. In addition, MEV extraction and optimization focus on maximizing the extracted value. In the following section, we explore how MEV simulation, extraction, and optimization methods are applied to better understand and manage MEV activities.
@@
| PROF [59] | Potential transaction exclusion in congestion scenarios, poses a centralization risk, and undermines transparency | ||||||
| CoWSwap | Poses a centralization risk and trade-off between efficiency and privacy | ||||||
@@ 8 MEV Simulation and Extraction Methods
In the context of MEV simulation, Qin et al. [18] proposed a method to simulate replacement front-running by duplicating a victim transaction, replacing only the sender’s address by the searcher’s address. The transaction is then executed locally on the highest block. If profitable, the searcher front-runs the victim transaction. A similar approach was developed by Stucke et al. [64] to simulate displacement front-running transactions272727https://github.com/zakstucke/ethereum-front-running, accessed on 25 February 2025. The approach executes the transactions locally on a Ganache fork to compute profitability. The approach also simulates sandwich transactions. This is by identifying victim transactions that perform swaps with a liquidity pool. It then front-runs the victim transaction by performing the same swap and back-runs it a reverse swap. Bogatyy and his team developed a simulator to perform displacement front-running transactions on the Bancor platform282828https://github.com/bogatyy/bancor, accessed on 05 September 2024. The simulator monitors the mempool for a victim transaction and executes a displacement transaction on the local Ethereum node. If profitable, the transaction is submitted to the network with a gas price higher than that of the victim via JSON RPC. The Subway bot292929https://github.com/libevm/subway, accessed on 05 September 2024 simulates sandwich transactions by monitoring the mempool for profitable opportunities. It determines optimal swap values and gas prices for front-running and back-running transactions, then executes them to capture profit.
@@
Regarding MEV extractions, Schwarz-Schilling et al. [65] and Öz et al. [66] discovered that the bid value submitted by a block builder in PBS increases with bid submission time. This increase is primarily driven by DEX-CEX arbitrage, where the risk associated with arbitrage decreases later in the slot as the DEX and CEX legs can be executed nearly simultaneously. As the time progresses within the block slot, the arbitrage profit becomes more certain, leading to an increase in the bid. Other non-primary factors for the increase might be the expansion of the public mempool and/or the submission of more private transaction bundles to the proposer. Consequently, [65, 66] suggested that an honest but rational block proposer might delay signing a block to optimize MEV. However, empirical analysis in these studies showed that proposers do not engage in such waiting games, likely due to concerns about potential reputation damage associated with self-serving behaviors. To optimize MEV, Babel and Baker [67] proposed MEV-peri, a peer selection approach that selects peers based on the value of transactions relayed by the peers and the latency of the received transactions. Empirical results reveal that peer selection using MEV-peri extracts more MEV compared to random and latency-based peer selection approaches.
@@
Bartoletti et al. [25] proposed a method to optimize MEV from dagwood sandwich transactions by modeling a transaction as a single-player single-round game. Similarly, Wang et al. [27] proposed a method to optimize MEV in the case of burger sandwich transactions. The proposed method involves a transaction selecting algorithm and an optimal attack algorithm. The selecting algorithm uses an iterative approach to determine a set of victim transactions that would yield optimal MEV. The optimal attack algorithm employs binary search to output parameters for an optimal front-running transaction based on the set of victim transactions. Babel et al. [63] introduced Lanturn303030https://github.com/lanturn-defi/lanturn, accessed on 25 February 2025, a learning-based framework for simulating MEV optimization strategies. It consists of an optimizer and a simulation environment. The optimizer, inspired by Genetic Algorithms, uses a bi-loop approach to develop a transaction sequence that maximizes profit. The outer loop determines the optimal order of transactions, while the inner loop fine-tunes transaction variables (e.g., token amounts in a swap transaction). The simulation environment evaluates the developed sequence by computing MEV at the current block height, executing transactions in a FIFO (First In, First Out) order on the local node. Experimental results showed that Lanturn extracts more MEV than Flashbots. Zhou et al. [68] introduced DEFIPOSER-ARB and DEFIPOSER-SMT to generate cyclic and non-cyclic arbitrage transactions for MEV extraction. DEFIPOSER-ARB identifies arbitrage opportunities across different DEXes by first constructing a graph of the current blockchain state and detecting cyclic transactions with potential profit. A local search then evaluates the profit from these opportunities, and the most profitable transaction is selected for inclusion in the next block. On the other hand, DEFIPOSER-SMT exploits non-cyclic profit-generating transactions by creating a state transition model of the DeFi system. The paths in the transition model represent potential profitable transaction sequences. Heuristics are then employed to prune the search space and execute the most profitable path. Li et al. [69] introduced a Role-Play strategy, wherein a searcher can simultaneously adopt multiple roles, such as trader, lender, borrower, and yield farmer, to optimize MEV exploitation across multiple DeFi applications. By mathematically modeling the interactions between these roles, the authors demonstrate how a searcher can extract significantly more value than traditional single-role exploits.
Table IV compares different MEV extraction methods based on the types of MEV transactions they consider, their inclusion of transaction fees, token transfer fees, wallet balance constraints, and liquidity pool reserves, as well as their evaluation approach, and optimization scope. Transaction fees refer to the gas fees paid by a searcher when executing MEV transactions. Overlooking these fees in the mathematical model for MEV optimization results in an overestimated profit, as the cost of performing transactions is not deducted from the computed gain. Furthermore, while performing transactions on DEXes, a fraction of the trading fee is paid to the liquidity provider. Accounting for this fee is essential to accurately compute MEV gains. It is also crucial to consider the searcher’s wallet balance to ensure that transactions can be successfully executed. Assuming an infinite token balance is unrealistic and does not reflect real-world constraints. Similarly, liquidity pool reserves must be taken into account to guarantee that both searcher and victim transactions can be executed successfully. The optimization scope is categorized as global if multiple MEV transactions, whether from the same or different strategies, are executed while accounting for continuous state changes after each MEV transaction. In contrast, local optimization considers only a single MEV strategy in isolation. However, a local optimization approach can still involve multiple transactions and multiple victims within that single strategy.
| MEV transactions | ||||||||||||||
| Extraction method | Front-running | Back-running | Sandwich | Arbitrage | Liquida-tion | Time-bandit | Transaction fee | Token transfer fee | Bound on wallet balance | Bound on liquidity pool reserve | Optimization | Evaluation | ||
| [67] | N/A | N/A | N/A | N/A | Local |
| [25] | Local | No execution | ||||||||||||
| [27] | Global |
| [63] | Global |
| [68] | N/A | N/A | N/A |
| [69] | N/A |
| : considered; : not considered; N/A: Not Applicable | ||||||||||||||